# Joint Identity Verification and Pose Alignment for Partial Fingerprints

### v3

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. In this paper, we propose a novel framework for joint identity verification and pose alignment of partial fingerprint pairs, aiming to leverage their inherent correlation to improve each other. To achieve this, we present a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset demonstrate that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
@@ I Introduction
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [10, 11]. While certain studies use neural networks to extract fixed-length region descriptors [12, 13, 14, 15], consequently reducing the dependence on point features, these approaches require either large effective area or exceedingly precise alignment. Considering this, some researchers simultaneously input fingerprint images and extract interrelated features to fully utilize complete information [7, 16], which effectively improves the matching performance on partial fingerprints. However, to the best of our knowledge, almost all existing fingerprint matching methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired regional features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [17, 18], in order to better balance robustness and speed in the overall process, rather than being made individually.
@@
In this paper, we propose a multi-task CNN-Transformer hybrid network which Jointly performs Identity verification and Pose alignment for partial fingerprints, called JIPNet. Paired images are input instead of their simplified features to provide more complete information, enabling more flexible and precise analysis. The intrinsic connection between feature correspondence and position correlation is exploited, aiming to promote each other to achieve higher performance. Naturally, the efficiency is also improved because these two tasks are compactly integrated. Inspired by works on natural images [19, 20, 21, 22], we utilize a shared weight convolution layers to extract regional features and transformers to capture both local and global correlations. Besides, a fingerprint enhancement pre-training task is specially designed to enhance the compatibility with different image modalities and various texture patterns.
@@ II-C Pairwise Interrelated Feature Based
Figure 3: An overview of JIPNet. Bars are presented on the left to indicate each phase, where the color definition refers to Fig. 2. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. * represents the number of channels are doubled after the corresponding convolution. Dotted parts are only used to assist training and could be pruned in practical verification tasks. The outputs ‘p’ and ‘H’ of respective task heads correspond to the classification probability (whether the input fingerprints come from the same finger) and rigid transformation parameters (relative translation and rotation) respectively.
@@ III-A Independent Feature Extraction
  • While blobs with random parameters which is added or subtracted to simulate fingerprint-specific noise in incomplete or contaminated scenarios [52].
@@
Figure 4: Illustration of fingerprint enhancement task for pre-training. The network architecture is shown on the left, where solid and dotted boxes indicate the pre-trained parameters will (or not) be loaded into corresponding modules in JIPNet. The right subfigure gives representative examples of inputs (original & augmented image) and targets (enhanced image) generation.
@@ IV Dataset Description
Comprehensive evaluations are conducted on extensive public datasets, including the representative public benchmarks NIST SD14 [23], FVC2002 DB1_A & DB3_A [24], FVC2004 DB1_A & DB2_A [25], and FVC2006 DB2_A [26]. In addition, we also use a private partial fingerprint dataset THU Small to further examine the performance in capacitive smartphone situations. Table I presents a detailed description about the composition and usage of these datasets. In particular, we merge several fingerprint databases into a larger dataset, called the Hybrid Database (Hybrid DB) , to make the implementation of various experiments more convenient and sufficient. On the other hand, we further evaluate the performance of cross-domain generalization on other separate datasets. Rolled fingerprint is utilized because it covers a more complete area at different pressing poses than plain fingerprints. Image examples of different datasets are shown in Fig. 6. It can be seen that datasets selected in our experiments contains diverse fingerprint impressions with various modalities (normal, dry, wet, polluted, wrinkled, etc.).
TABLE I: All Fingerprint Datasets Used in Experiments.
@@ IV-B Matching Protocols
In order to balance the number of genuine / impostor matches and different fingerprint types / modalities, we implemented specific matching protocols during training and testing. For the FVC series, each fingerprint is combined with all mated fingerprints (with order) as genuine match pairs, and the first image of each pair is combined with a random impression of another finger as a new impostor match pair. Taking FVC2002 DB1_A as an example, a total of $8\times 7\times 100\times 2=11,200$ image pairs are combined in this way. This protocol is also applied to NIST SD14 with a slightly different that symmetric matches are avoided, considering the large number of rolled images. All combined pairs of full fingerprints are then randomly cropped to simulate several partial fingerprint scenes, with the number of executions set to $4$ (NIST SD14, FVC2004 DB1_A & DB2_A, FVC2006 DB2_A) for training  /  testing and $1$ (FVC2002 DB1_A & DB3_A) for testing. As mentioned above, some representative datasets are merged as Hybrid DB and subsequently divided, secifically $430,768$ for training and $22,672$ for testing, which are distinguished by suffixes ‘_A’ and ‘_B’ in the following experiments. The selection of genuine matches in real partial fingerprint dataset THU Small also follows this principle, while pairing the first impression of every finger with each other as impostor pairs. All detailed usage information is listed in corresponding footnotes of Table I.
@@ V Experiments
In particular, since the effective information of partial fingerprints is significantly reduced compared to full-size fingerprints, the original pose rectification pre-step adopted in DeepPrint [12] and DesNet [13] cannot be performed as expected. Therefore, in this paper we use AlignNet, part of PFVNet [7], to estimate and correct the relative pose of image pairs as substitute, which is marked by subscript ‘*’.
@@ V-B Score Distribution
Moreover, three common indicators are further computed and reported in Table II to quantitatively measure the differences between genuine and impostor probability distributions. Among them, the Jensen–Shannon (JS) divergence is defined as:
@@
TABLE II: Probability Distribution Distance Between Genuine and Impostor Scores on Hybrid DB_B with Image Size $160\times 160$.
Method JS EMD MMD
A-KAZE [6] 0.42 0.20 0.72
VeriFinger [9] 0.70 0.59 3.21
DeepPrint* [12] 0.59 0.36 2.03
DesNet* [13] 0.47 0.17 1.10
PFVNet [7] 0.73 0.43 3.01
[…]
According to the results in Fig. 7 and Table II, our method outperforms all other methods, while VeriFinger [9] and PFVNet [7] each have sub-optimal performance in some parts. In particular, the distinguishing ability of DesNet* [13] is not ideal because it is designed based on the prior assumption of high-precision alignment, which is difficult to satisfy in current partial fingerprint situations.
@@ V-C Matching Performance
TABLE III: Matching Performance of State-of-the-Art Algorithms and Our Approach on Four Representative Datasets of Partial Fingerprint with Three Different Sizes.
@@
TABLE IV: Matching Performance of Different Score Level Fusion Strategies with image size $160\times 160$.
@@
[…]
Table III presents the corresponding matching performance of state-of-the-art algorithms and our proposed approach on several representative datasets of fingerprints with different sizes. Furthermore, we group the genuine matches on Hybrid DB_B according to their relative rigid transformation relationship and compare these algorithms under different rotations and translations using EER. Fig. 8 shows the corresponding comparisons. Additionally, the Detection Error Tradeoff (DET) curves on all test databases are calculated in Fig. .1 to provide more complete information for analysis.
These encouraging experimental results strongly demonstrates the superiority of our proposed JIPNet, which outperforms previous advanced methods in almost all cases. PFVNet [7] also shows remarkable performance, especially in indicators of ACC and AUC. However, it can be observed that the relative performance deteriorates markedly when training and testing data cross domains. Region representation based methods[12, 13] are inferior to others, which is reasonable because the highly generalized form of fixed-length descriptors sacrifices many potential details for faster search speed. Specifically, although localized descriptors with deeper dimensions are designed in DesNet [13], their effectiveness is limited, as stated in Section V-B. On the other hand, the performance of all algorithms drops significantly as the image size decreases or distance increases, but our scheme still leads and exhibits attractive robustness and stability. It is worth noting that the impact of rotation is not evident in Fig. 8, because key points based matching score do not require pose rectification and the prediction error in other algorithms remains at a similar level across different relative angles (confirmed in the following experiments).
As analyzed above, these compared methods focus on different aspects of fingerprint attributes, exhibiting differentiated trends in score distribution and matching advantages. This phenomenon motivates us to conduct multiple fusion strategies to further observe their complementarity. Specifically, scores from two sets are linearly weighted and summed with the best classification accuracy. As represented in Table IV, most fusions effectively improves the corresponding performance, especially the introduction of minutiae information (VeriFinger) into other methods that rely on texture features, while our method still leads the way.
@@ V-D Alignment Accuracy
V-E Visual Analysis
@@ V-F Efficiency
V-G Ablation Study
TABLE V: Ablation Study of the Proposed Network with Different Modules and Strategies on FVC2002 DB1_A with image size $160\times 160$.
Encoder * Decoder * + Align Head * EER
CNN ViT CNN ViT
0.25
0.21
0.47
0.29
0.17
0.07
[…]
  • §$\S$
    indicates the corresponding module has been pretrained.
    each represents specific stages in proposed network, specifically Encoder for Stage 1 & 2, Decoder for Stage 3 & 4 - C, Align Head for Stage 4 - R.
Extensive ablation experiments are conducted to examine the effects of different modules and strategies introduced in our network. As shown in Table V, overall, the CNN based approach outperforms Vision transformer (ViT) based approach, while the CNN-Transformer hybrid series achieves the best performance. We attribute it to the respective advantages of convolutional and transformer blocks in extracting local features and establishing global connections. In addition, the introduction of parallel relative alignment estimation head achieves great success in all three types of network architectures, which proves that there is indeed an exploitable promotion relationship between pose prediction and identity verification. At the same time, the additional alignment head can be pruned during deployment so that the matching speed will not be affected. Finally, the comparison between last two rows demonstrated that our proposed lightweight pre-training task can further improve the performance at a small extra training time cost.

### v4

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. In this paper, we propose a novel framework for joint identity verification and pose alignment of partial fingerprint pairs, aiming to leverage their inherent correlation to improve each other. To achieve this, we present a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset demonstrate that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods. Code is available at: https://github.com/XiongjunGuan/JIPNet.
@@ I Introduction
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [10, 11]. While certain studies use neural networks to extract fixed-length region descriptors [12, 13, 14, 15], consequently reducing the dependence on point features, these approaches require either large effective area or exceedingly precise alignment. Considering this, some researchers simultaneously input fingerprint images and extract interrelated features to fully utilize complete information [7, 16], which effectively improves the matching performance on partial fingerprints. However, to the best of our knowledge, almost all existing fingerprint matching methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired regional features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where decision-makers need to consider whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [17, 18]. It can be considered as a mixed process, as the relative alignment relationship is confirmed (one of the eight lateral positions for a regular puzzle) simultaneously when two pieces are determined to be adjacent. On the other hand, an input piece can be classified as outside (mismatched) when it cannot be properly connected to the reference piece at any position. The comprehensive consideration of identity verification and pose alignment can be transferred to fingerprints, which is feasible because previous works [7, 19] have demonstrated the complementarity and compatibility between above two tasks. In addition, the paradigm of joint decision can better balance the robustness and speed of overall process, rather than performing them individually.
@@
In this paper, we propose a multi-task CNN-Transformer hybrid network which Jointly performs Identity verification and Pose alignment for partial fingerprints, called JIPNet. Paired images are input instead of their simplified features to provide more complete information, enabling more flexible and precise analysis. The intrinsic connection between feature correspondence and position correlation is exploited, aiming to promote each other to achieve higher performance. Unlike previous independent two-stage approaches, the estimation of identity and pose are jointly executed in our proposed framework to leverage their complementarity. Naturally, the efficiency is also improved because these two tasks are compactly integrated. Inspired by works on natural images [20, 21, 22, 23], we utilize a shared weight convolution layers to extract regional features and transformers to capture both local and global correlations. Besides, a fingerprint enhancement pre-training task is specially designed to enhance the compatibility with different image modalities and various texture patterns.
@@ II-C Pairwise Interrelated Feature Based
Figure 3: An overview of JIPNet. Paired fingerprint patches with the same shape are input, specifically $160\times 160$, $120\times 120$, or $96\times 96$ in this paper. The outputs ‘p’ and ‘H’ of respective task heads correspond to the classification probability (whether the input fingerprints come from the same finger) and rigid transformation parameters (relative translation and rotation) respectively. Detailed structure is shown in Table I. Bars are presented on the left to indicate each phase, where the color definition refers to Fig. 2. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. * represents the number of channels are doubled after the corresponding convolution. Dotted parts are only used to assist training and could be pruned in practical verification tasks.
TABLE I: Detailed configurations of JIPNet. A pair of $160\times 160$ fingerprint patches are inputted as an example.
Phase Layer Operator Output shape
Independent Feature Extraction Stem $3\times 3$ Conv $32,160\times 160$
Stage 1 Encoder Block $64,80\times 80$
Encoder Block $128,40\times 40$
Encoder Block $256,20\times 20$
Stage 2 Residual Block ×$\times 4$ $264,20\times 20$
Interrelated Feature Extraction Positional Encoding $264,400$
Stage 3 Transformer Block ×$\times 8$ $264,400$
Patch Merging $384,100$
Positional Encoding $384,100$
Identity Verification Stage 4 - C Transformer Block ×$\times 4$ $384,100$
Concatenate $768,100$
Classifier Linear Projection $1,1$
Sigmoid $1,1$
Pose Prediction Stage 4 - R Transformer Block ×$\times 4$ $384,100$
Concatenate $768,100$
Regressor Linear Projection $4,1$
@@ III-A Independent Feature Extraction
Figure 4: Illustration of fingerprint enhancement task for pre-training. The network architecture is shown on the left, where solid and dotted boxes indicate the pre-trained parameters will (or not) be loaded into corresponding modules in JIPNet. Detailed structure is shown in Table II. The right subfigure gives representative examples of inputs (original & augmented image) and targets (enhanced image) generation.
TABLE II: Detailed configurations of the fingerprint enhancement network in pre-training task. A pair of $160\times 160$ fingerprint patches are inputted as an example.
Layer Operator Output shape
Stem $3\times 3$ Conv $32,160\times 160$
Stage 1 Encoder Block $64,80\times 80$
Encoder Block $128,40\times 40$
Encoder Block $256,20\times 20$
Stage 2 FE Block ×$\times 2$ $256,20\times 20$
Stage 3 Decoder Block $128,40\times 40$
Decoder Block $64,80\times 80$
Decoder Block $32,160\times 160$
Seg. Head $3\times 3$ Conv $1,160\times 160$
Sigmoid $1,160\times 160$
@@
  • White blobs with random parameters which is added or subtracted to simulate fingerprint-specific noise in incomplete or contaminated scenarios [53].
@@ IV Dataset Description
Comprehensive evaluations are conducted on extensive public datasets, including the representative public benchmarks NIST SD14 [24], FVC2002 DB1_A & DB3_A [25], FVC2004 DB1_A & DB2_A [26], and FVC2006 DB2_A [27]. In addition, we also use a private partial fingerprint dataset THU Small to further examine the performance in capacitive smartphone situations. Table III presents a detailed description about the composition and usage of these datasets. In particular, we merge several fingerprint databases into a larger dataset, called the Hybrid Database (Hybrid DB) , to make the implementation of various experiments more convenient and sufficient. On the other hand, we further evaluate the performance of cross-domain generalization on other separate datasets. Rolled fingerprint is utilized because it covers a more complete area at different pressing poses than plain fingerprints. Image examples of different datasets are shown in Fig. 6. It can be seen that datasets selected in our experiments contains diverse fingerprint impressions with various modalities (normal, dry, wet, polluted, wrinkled, etc.).
TABLE III: All Fingerprint Datasets Used in Experiments.
@@ IV-B Matching Protocols
In order to balance the number of genuine / impostor matches and different fingerprint types / modalities, we implemented specific matching protocols during training and testing. For the FVC series, each fingerprint is combined with all mated fingerprints (with order) as genuine match pairs, and the first image of each pair is combined with a random impression of another finger as a new impostor match pair. Taking FVC2002 DB1_A as an example, a total of $8\times 7\times 100\times 2=11,200$ image pairs are combined in this way. This protocol is also applied to NIST SD14 with a slightly different that symmetric matches are avoided, considering the large number of rolled images. All combined pairs of full fingerprints are then randomly cropped to simulate several partial fingerprint scenes, with the number of executions set to $4$ (NIST SD14, FVC2004 DB1_A & DB2_A, FVC2006 DB2_A) for training  /  testing and $1$ (FVC2002 DB1_A & DB3_A) for testing. As mentioned above, some representative datasets are merged as Hybrid DB and subsequently divided, secifically $430,768$ for training and $22,672$ for testing, which are distinguished by suffixes ‘_A’ and ‘_B’ in the following experiments. The selection of genuine matches in real partial fingerprint dataset THU Small also follows this principle, while pairing the first impression of every finger with each other as impostor pairs. All detailed usage information is listed in corresponding footnotes of Table III.
@@ V Experiments
  • AFR-Net [15], an attention-driven fingerprint recognition network that extracts complementary global representations through ViT and ResNet embeddings.
@@
In particular, since the effective information of partial fingerprints is significantly reduced compared to full-size fingerprints, the original pose rectification pre-step adopted in DeepPrint [12], DesNet [13] and AFR-Net [15] cannot be performed as expected. Therefore, in this paper we use AlignNet, part of PFVNet [7], to estimate and correct the relative pose of image pairs as substitute, which is marked by subscript ‘*’.
@@ V-B Score Distribution
TABLE IV: Probability Distribution Distance Between Genuine and Impostor Scores on Hybrid DB_B with Image Size $160\times 160$.
Method JS EMD MMD
A-KAZE [6] 0.42 0.20 0.72
VeriFinger [9] 0.70 0.59 3.21
DeepPrint* [12] 0.59 0.36 2.03
DesNet* [13] 0.47 0.17 1.10
AFR-Net* [15] 0.26 0.15 0.43
PFVNet [7] 0.73 0.43 3.01
[…]
Moreover, three common indicators are further computed and reported in Table IV to quantitatively measure the differences between genuine and impostor probability distributions. Among them, the Jensen–Shannon (JS) divergence is defined as:
@@
According to the results in Fig. 7 and Table IV, our method outperforms all other methods, while VeriFinger [9] and PFVNet [7] each have sub-optimal performance in some parts. In particular, the distinguishing ability of DesNet* [13] is not ideal because it is designed based on the prior assumption of high-precision alignment, which is difficult to satisfy in current partial fingerprint situations. Another noteworthy phenomenon is that the discriminative ability of AFR-Net [15] is not satisfactory in this small image patch scenario, despite its advanced performance on rolled/plain fingerprint matching. We believe that its attention driven architecture effectively increases the ability to express local textures. However, this high sensitivity may actually limit robustness when there is limited available information (small effective fingerprint area) or excessive interference (limited overlap ratio).
@@ V-C Matching Performance
TABLE V: Matching Performance of State-of-the-Art Algorithms and Our Approach on Four Representative Datasets of Partial Fingerprint with Three Different Sizes.
@@
Figure 9: DET curves of state-of-the-art algorithms and our approach on partial fingerprints of $160\times 160$. Solid and dotted lines represent deep learning methods and traditional methods respectively.
[…]
Table V presents the corresponding matching performance of state-of-the-art algorithms and our proposed approach on several representative datasets of fingerprints with different sizes. Furthermore, we group the genuine matches on Hybrid DB_B according to their relative rigid transformation relationship and compare these algorithms under different rotations and translations using EER. Fig. 8 shows the corresponding comparisons of representative methods for each type of matching framework. Additionally, the Detection Error Tradeoff (DET) curves on all test databases are calculated in Fig. 9 to provide more complete information for analysis.
These encouraging experimental results strongly demonstrates the superiority of our proposed JIPNet, which outperforms previous advanced methods in almost all cases. PFVNet [7] also shows remarkable performance, especially in indicators of ACC and AUC. However, it can be observed that the relative performance deteriorates markedly when training and testing data cross domains. Region representation based methods[12, 13, 15] are inferior to others, which is reasonable because the highly generalized form of fixed-length descriptors sacrifices many potential details for faster search speed. On the other hand, emphasizing certain texture features located in non overlapping areas may also lead to incorrect mismatches. That is to say, genuine matches with very small overlapping areas and impostor matches with very similar local patterns lack sufficient stable discrimination under such regional representations, which is also reflected in Fig. 7. Specifically, although localized descriptors with deeper dimensions are designed in DesNet [13], their effectiveness is limited, as stated in Section V-B. On the other hand, the performance of all algorithms drops significantly as the image size decreases or distance increases, but our scheme still leads and exhibits attractive robustness and stability. It is worth noting that the impact of rotation is not evident in Fig. 8, because key points based matching score do not require pose rectification and the prediction error in other algorithms remains at a similar level across different relative angles (confirmed in the following experiments).
TABLE VI: Matching Performance of Different Score Level Fusion Strategies with image size $160\times 160$.
@@
As analyzed above, these compared methods focus on different aspects of fingerprint attributes, exhibiting differentiated trends in score distribution and matching advantages. This phenomenon motivates us to conduct multiple fusion strategies to further observe their complementarity. Specifically, scores from two sets are linearly weighted and summed with the best classification accuracy. As represented in Table VI, most fusions effectively improves the corresponding performance, especially the introduction of minutiae information (VeriFinger) into other methods that rely on texture features, while our method still leads the way.
@@ V-D Alignment Accuracy
V-E Ablation Study
Extensive ablation experiments are conducted to examine the effects of different modules and strategies introduced in our network. EER and $\mathcal{L}_{\mathrm{reg}}$ (defined in Equation 9) are used to measure the performance of matching and alignment, respectively. As shown in Table VII, overall, the CNN based approach outperforms Vision transformer (ViT) based approach, while the CNN-Transformer hybrid series achieves the best performance. We attribute it to the respective advantages of convolutional and transformer blocks in extracting local features and establishing global connections. In addition, the introduction of parallel relative alignment estimation head achieves great success in EER metric across all three types of network architectures, which proves that there is indeed an exploitable promotion relationship between pose prediction and identity verification. Furthermore, the joint estimation strategy also shows certain positive effect on pose prediction task when applied in networks containing ViT decoders. At the same time, the additional alignment head can be pruned during deployment so that the matching speed will not be affected. Finally, the comparison between last two rows demonstrated that our proposed lightweight pre-training task can further improve the performance at a small extra training time cost. A convincing explanation is that it ensures the extracted features contain complete pattern information of fingerprint ridges, while also performing a certain degree of denoising. In this way, the inputs for subsequent two modules, namely identity verification and pose prediction, is transferred from the image level to the feature level, which assists the network to better focus on key features and perform joint optimization more seamlessly.
TABLE VII: Ablation Study of the Proposed Network with Different Modules and Strategies on FVC2002 DB1_A with image size $160\times 160$.
[…]
CNN ViT CNN ViT
0.25 -
- 18.2
0.21 20.1
0.47 -
- 27.2
0.29 26.7
0.17 -
- 10.6
0.07 9.2
[…]
  • §$\S$
    indicates the corresponding module has been pretrained through the process shown in Figure 4.
    each represents specific stages in proposed network, specifically Encoder for Stage 1 & 2, Decoder for Stage 3, ID Head for Stage 4 - C, Pose Head for Stage 4 - R.
V-F Visual Analysis